Suppose there is a country called Alisa-opolis. In this country, we have the following information:

Number of people 16 and over: 1,900,000

Number of people in institutions like prison: 75,000

Number of people in the military: 225,000

Number of people who are working: 900,000

Number of people who are unemployed: 100,000

What is the unemployment rate in this country?

Group of answer choices

5.5%

10.0%

12.5%

Not enough information

Using the information from the previous problem, what is the Labor Force Participation rate for this country?

Group of answer choices

50.3%

62.5%

77%

85.3%
Answer #1

1. 10%

Explanation:

Unemployment rate = unemployed/labor force * 100 = unemployed / (employed + unemployed) * 100 = 100,000/ (900,000 + 100,000)*100 = 100,000/1,000,000 * 100 = 1/10 * 100 = 10%

2. 62.5%

Explanation:

Labor force participation rate = labor force/noninstitutionalized, civilian working-age population* 100

Here, noninstitutionalized, civilian working-age population = Number of people 16 and over - (people in military + people in prison) = 1,900,000 - (225,000 + 75,000) = 1,900,000 - 300,000 = 1,600,000

Labor force = employed + unemployed = 900,000 + 100,000 = 1,000,000.

So, the labor force participation rate = labor force/noninstitutionalized, civilian working-age population* 100 = 1,000,000/1,600,000 * 100 = 0.625 * 100 = 62.5%